OpenStudy (anonymous):

what is the simplified form of (y^2-y-6 / y^2-4)?

OpenStudy (anonymous):

factor and cancel the numerator factors as \(y^2-y-6=(y+2)(y-3)\)

OpenStudy (anonymous):

(You will want to factor the denominator as well) Thankfully, it is a difference of squares.

OpenStudy (anonymous):

so if I got (y-3/y-2) how would I find the restriction?

OpenStudy (anonymous):

Your restrictions derive from the initial expression's denominator. So, going back to your initial (factored) expression of \[\frac{ \left( y-3 \right)\left( y+2 \right) }{ \left( y-2 \right)\left( y+2 \right) }\] you find the restrictions by setting your denominator equal to zero and solving.

OpenStudy (anonymous):

so would it be positive 2 or y ≠ ±2

OpenStudy (anonymous):

Nice! It would be \[y \neq \pm 2\]

OpenStudy (anonymous):

thanks! helped alot :)

OpenStudy (anonymous):

Of course! Have a good day. :D